What volume (in mL) of 0.4050M calcium chloride reacts completely with 25.00mL of 0.2800M silver nitrate?


1
Expert's answer
2021-06-25T06:52:35-0400

"\u0421(A"gNO3)*V(AgNO3)= C(CaCl2) *V (CaCl2)

V(CaCl2)= C (AgNO3)*V (AgNO3)/ C(CaCl2)= 25*0.28/0.405=17.28 ml